# Enhancing Oscillator-Based Ising Machine Models with Amplitude Dynamics and Polynomial Interactions

### v1

@@ IV.4 Binarization Behavior: SAT Problems
Figure 4: Example phase plots for a non-binarized (left), poorly binarized (middle) and a well-binarized (right) system with axis of binarization $\vec{b}$.
Similar to the quadratic case, there are multiple choices to formulate a cubic oscillator model. Our empirical analysis on 3-SAT problems shows that the model defined by Eq. (15) attains the highest solution quality, but only weakly binarizes. Conversely, replacing the terms in Eq. (15) with the complete form in Eq. (22) yields a new energy function which strongly binarizes but suffers from inferior optimization performance. In our experiments, this variant solved 61.8%, 17%, 2%, and 0% of the instances for the uf-50, uf-75, uf-100, and uf-150 benchmarks, respectively, based on 100 independent trials per instance, which is similar to the baseline we compared with. This trade-off motivates the question: Can we design an oscillator network that simultaneously (i) achieves high solution quality, (ii) respects the Ising Hamiltonian, and (iii) yields binary spin values at convergence?
To address this, we examine the contributions of each term in the formulation Eq. (15) independently. Linear and cubic terms inherently attain weak binarization, as their Hessians have non-zero curvature along $(1)^{T},(1,1,1)^{T}$. The primary difference between the Hamiltonians lies with the quadratic terms. As we showed in the previous subsection, C-potential quadratic Hamiltonians are strongly convex in the vicinity of $\{0,\pi\}$ phase points, promoting strong binarization. In contrast, the 1-potential does not binarize, as the system is invariant to global phase shifts.
Figure 6: Cumulative distribution of final energy values for the different binarization strategies discussed.
Therefore, introducing a time-varying quadratic Wirtinger potential can provide strong binarization force while preserving the optimization performance of the oscillator network. More specifically, we can get a model with stronger binarization tendency by modifying the quadratic terms (order $k=2$) to be the complete form we previously defined in Eq. (22) as follows:
@@
$\quad+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*}\Bigr{)}$
@@
It is important to note that after binarization, both Hamiltonian (15) and Eq. (35), as well as any combination of them, yield identical values to the discrete Ising formulation, although binarization introduces a slight modification in the energy landscape. To demonstrate the effectiveness of our binarization strategies, we evaluated four scenarios on five uf-100 3-SAT instances:
  • None: The system evolves under Hamiltonian (15).
  • Static Potential: The system evolves under Hamiltonian (35).
  • Annealed SHIL: The system uses Hamiltonian (15) augmented with a SHIL signal, where the coefficient increases linearly.
  • Annealed Potential: The system starts with Hamiltonian (15) and linearly transitions to Hamiltonian (35) to achieve binarization.
Fig. 5 presents the distribution of the final oscillator phases. In the None scenario, the distributions are weakly peaked at 00 and $\pi$. However, a significant proportion of the phases are spread between the peaks, indicating little inherent binarization. In contrast, both the SHIL and Annealed strategies exhibit a strong binarization tendency.
The cumulative distribution plot of the final solution energy in Fig. 6 illustrates the overall performance (solution quality). The narrower figure of merit (FoM) should be just the probability of reaching the ground state (energy=0). However, we note that these are stochastic results and one way to better see through the underlying reality is to see the distribution of all solutions. (Another way is to show the narrow FoM over more test benchmarks.) In the cumulative distribution, the higher the curve the better. As we can see, the SHIL and Annealed strategies deliver performance that is similar to or slightly better than that achieved by the system evolving under Hamiltonian (15) alone. We hypothesize that the induced binarization acts as a form of stochastic perturbation, which may help the system escape local minima during the optimization process.
V DISCUSSION
Phase reduction approaches, including the Kuramoto model, fail to capture the full dynamics of oscillator-based Ising machines in real-world applications. We present a comprehensive Hopf oscillator model that explicitly incorporates amplitude dynamics, addressing this fundamental limitation. Our model demonstrates superior performance, consistently converging to lower energy states compared to phase-only models.
Our investigation focuses on model-level improvements, specifically examining the energy distribution when solving 3-SAT problems and quantifying the number of solvable problems without heuristic enhancements. While recent work [24, 57, 72] has shown that incorporating heuristic mechanisms can significantly boost solver performance, our fundamental model improvements provide a more robust foundation for such enhancements.
Incorporating amplitude dynamics introduces an additional degree of freedom. The use of a complex oscillator representation paves the way for further research. We interpret the local dynamics term $f(z_{i})$ in the oscillator model (Eq. (12)) not only as a mechanism to preserve physical oscillator properties but also as a means to improve the performance of the oscillator-based Ising machine. Here, we provide a perspective for future studies: oscillator-based Ising machines rely on subharmonic injection locking (SHIL) and nonlinearity to achieve better performance on MAX-CUT problems [28, 61], whereas Linear Ising Machines (LIMs) like BRIM [52] do not. By engineering the local dynamics, one can steer the oscillator-based Ising machine to mimic the dynamics of LIMs. For example, in quadratic MAX-CUT problems, the oscillator dynamics $f(z_{i})$ could be designed to preserve the real part of the oscillator representation $z_{i}$ while adjusting the imaginary part to maintain essential oscillator properties. This approach shifts the optimization focus from hard-to-implement system-level strategies to tuning inherent oscillator properties, although further experimental validation remains necessary.
VI CONCLUSION
Our work is driven by the need for a more accurate model for oscillator-based Ising machines. We observed several fundamental inconsistencies in existing phase-amplitude models. By addressing these issues, we propose a model that incorporates both phase and amplitude dynamics, features a physically meaningful Lyapunov function, and remains compatible with the widely used phase-only models in oscillator-based Ising machine implementations. Experiments show that our model outperforms existing phase-amplitude approaches in optimization performance and exhibits a correct energy minimization tendency. Furthermore, we propose a method to generalize the model for constructing higher-order Ising machines and an implementable approach for extracting solutions from the oscillators. These findings contribute to both the theoretical understanding and practical realization of oscillator-based Ising machines, offering valuable directions for future research and development in this promising field.

### v2

@@ IV.4 Binarization Behavior: SAT Problems
Similar to the quadratic case, there are multiple choices to formulate a cubic oscillator model. Our empirical analysis on 3-SAT problems shows that the model defined by Eq. 15 attains the highest solution quality, but only weakly binarizes. Conversely, replacing the terms in  Eq. 15 with the complete form in Eq. 24 yields a new energy function which strongly binarizes but suffers from inferior optimization performance. In our experiments, this variant solved 61.8%, 17%, 2%, and 0% of the instances for the uf-50, uf-75, uf-100, and uf-150 benchmarks, respectively, based on 100 independent trials per instance, which is similar to the baseline we compared with. This trade-off motivates the question: Can we design an oscillator network that simultaneously (i) achieves high solution quality, (ii) respects the Ising Hamiltonian, and (iii) yields binary spin values at convergence?
To address this, we examine the contributions of each term in the formulation Eq. 15 independently. Linear and cubic terms inherently attain weak binarization, as their Hessians have non-zero curvature along $(1)^{\top},(1,1,1)^{\top}$. The primary difference between the Hamiltonians lies with the quadratic terms. As we showed in the previous subsection, C-potential quadratic Hamiltonians are strongly convex in the vicinity of $\{0,\pi\}$ phase points, promoting strong binarization. In contrast, the 1-potential does not binarize, as the system is invariant to global phase shifts.
Therefore, introducing a time-varying quadratic Wirtinger potential can provide a strong binarization force while preserving the optimization performance of the oscillator network. More specifically, we can get a model with a stronger binarization tendency by selecting $\mathcal{H}^{C}_{2}$ for the quadratic terms as follows:
@@
$\quad+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*}\Bigr{)}+\text{constant}.$
It is important to note that after binarization, both Hamiltonian Eq. 15 and Eq. 34, as well as any combination of them, yield identical values to the discrete Ising formulation, although binarization introduces a slight modification in the energy landscape. To demonstrate the effectiveness of our binarization strategies, we evaluated four scenarios on 45 uf100-430 3-SAT instances:
  • None: The system evolves under Hamiltonian Eq. 15.
  • Static Potential: The system evolves under Hamiltonian Eq. 34.
  • Annealed SHIL: The system uses Hamiltonian Eq. 15 augmented with a SHIL signal, where the coefficient increases linearly.
  • Annealed Potential: The system starts with Hamiltonian Eq. 15 and linearly transitions to Hamiltonian Eq. 34 to achieve binarization.
Fig. 4 presents the distribution of the final oscillator phases. In the None scenario (a), the distributions are weakly peaked at 0 and $\pi$. However, a significant proportion of the phases are spread between the peaks, indicating little inherent binarization. In contrast, the SHIL, Static, and Annealed strategies (b)-(d) exhibit a strong binarization tendency.
Figure 5: Cumulative distribution of final energy values for the different binarization strategies discussed. Each line represents the distribution of 45 problems from uf100-430 with 100 trials each. 99% confidence intervals are provided as shaded lines around each estimate, however they are typically too narrow to be visible.
Fig. 5 shows the cumulative final energy distribution of each binarization strategy. The distributions represent 45 problems from uf100-430 ($n=100$, $m=430$), each with 100 runs. We use bootstrap resampling with $10^{4}$ resamples to obtain 99% confidence intervals (shaded area around lines); however, in most cases, they are too narrow to be easily visible.
In lieu of final energy, a narrower figure of merit (FoM) should be just the probability of reaching the ground state (energy=0). However, we note that these are stochastic results, and one way to better see through the underlying reality is to see the distribution of all solutions. (Another way is to show the narrow FoM over more test benchmarks.) In the cumulative distribution, the higher the curve, the better. As we can see, the SHIL and Annealed strategies deliver performance that is similar to or slightly better than that achieved by the system evolving under Hamiltonian Eq. 15 alone. We hypothesize that the induced binarization acts as a form of stochastic perturbation, which may help the system escape local minima during the optimization process.
V DISCUSSION
Phase reduction approaches, including the Kuramoto model, fail to capture the full dynamics of oscillator-based Ising machines in real-world applications. We present a comprehensive Hopf oscillator model that explicitly incorporates amplitude dynamics, addressing this fundamental limitation. Compared to the existing baseline, our model defines a clear energy function with corresponding dynamics and achieves better performance.
Our study focuses on the theoretical soundness of phase–amplitude models for oscillator-based Ising machines. Recent studies [52, 89, 105] demonstrate that integrating heuristic mechanisms substantially improve Ising-machine-based 3-SAT solver performance. Our model enhancements establish a solid foundation for these advances.
Incorporating amplitude dynamics introduces an additional degree of freedom. The use of a complex oscillator representation paves another way for further research. We interpret the local dynamics term $f(z_{i})$ in the oscillator model (Eq. 12) not only as a mechanism to preserve physical oscillator properties but also as a means to improve the performance of the oscillator-based Ising machine. Here, we provide a perspective for future studies: oscillator-based Ising machines rely on binarization and nonlinearity to achieve better performance on MAX-CUT problems [56, 93], whereas Linear Ising Machines (LIMs) like BRIM [35] do not. By engineering the local dynamics, one can steer the oscillator-based Ising machine to mimic the dynamics of LIMs. For example, in quadratic MAX-CUT problems, the oscillator dynamics $f(z_{i})$ could be designed to preserve the real part of the oscillator representation $z_{i}$ while adjusting the imaginary part to maintain essential oscillator properties. This approach shifts the optimization focus from system-level strategies to tuning inherent oscillator properties, although further experimental validation remains necessary.
VI CONCLUSION
Our work is driven by the need for a more accurate model for oscillator-based Ising machines. We observed several fundamental inconsistencies in existing phase-amplitude models. By addressing these issues, we propose a model that incorporates both phase and amplitude dynamics, features a physically meaningful Lyapunov function, and remains compatible with the widely used phase-only models in oscillator-based Ising machine implementations. Experiments show that our model outperforms existing phase-amplitude approaches in optimization performance and exhibits a correct energy minimization tendency.
Our phase-amplitude model provides a more faithful framework for modeling oscillator-based Ising machines when amplitude dynamics cannot be neglected. This includes spintronic oscillators that naturally exhibit coupled amplitude and phase dynamics, NEMS oscillators operating in regimes where amplitude variations are significant, and emerging photonic systems. By providing a rigorous mathematical foundation that captures these dynamics, our model enables more accurate simulations and predictions of system behavior, which is crucial for the design and optimization of OIMs. Accurate modeling further enhances practical applications, including portfolio optimization, drug discovery, wireless communication, and a wide range of other combinatorial optimization problems.
Furthermore, we propose a method to generalize the model for constructing higher-order Ising machines and an implementable approach for extracting solutions from the oscillators. These findings contribute to both the theoretical understanding and practical realization of oscillator-based Ising machines, offering valuable directions for future research and development in this promising field.
ACKNOWLEDGMENTS
This work was supported in part by NSF under Awards No. 2231036 and No. 2233378, and by DARPA under contract No. FA8650-23-C-7312.
DATA AVAILABILITY
The data that support the findings of this article are openly available [106].
APPENDIX
Appendix A Example: Mapping a 3-SAT problem to Ising Hamiltonian
We illustrate the mapping on the single clause
$C=(x_{1}\lor\neg x_{2}\lor x_{3}).$
Our goal is to construct a Hamiltonian $H_{C}$ for this clause. When $H_{C}$ is minimized to 0, the clause is satisfied; otherwise, it contributes 1 to the total energy.
Mapping Method: Each Boolean $x_{i}\in\{0,1\}$ is represented by an Ising spin $s_{i}\in\{+1,-1\}$, with
$s_{i}=+1\iff x_{i}=\text{true},\quad s_{i}=-1\iff x_{i}=\text{false}.$
Step 1: Literal energy terms. For each literal $\ell$, define
$f(\ell)=\begin{cases}1-s_{i},&\ell=x_{i},\\ 1+s_{i},&\ell=\neg x_{i}.\end{cases}$
In our clause,
$f(x_{1})=1-s_{1},\quad f(\neg x_{2})=1+s_{2},\quad f(x_{3})=1-s_{3},$
according to the mapping method, $f(\ell)=0$ when the literal is true, and $f(\ell)=2$ otherwise.
Step 2: Clause energy by multiplication. To ensure the energy is nonzero only when all three literals fail, we take the product:
@@ Appendix A Example: Mapping a 3-SAT problem to Ising Hamiltonian
Step 3: Normalization. Dividing by 8 yields the clause Hamiltonian
$H_{C}=\frac{(1-s_{1})\,(1+s_{2})\,(1-s_{3})}{8}=\begin{cases}0,&\text{clause satisfied},\\ 1,&\text{clause unsatisfied}.\end{cases}$
Finally, for a full 3-SAT problem with clauses $C_{1},\dots,C_{M}$, we sum these terms:
$\mathcal{H}=\sum_{j=1}^{M}H_{C_{j}},$
so $\mathcal{H}$ counts exactly the number of unsatisfied clauses.
Appendix B Wirtinger Calculus for Non-Holomorphic Hamiltonians
Let $z=x+iy$ be a complex $n$-dimensional vector, where $x,y\in\mathbb{R}^{n}$. When the Hamiltonian $\mathcal{H}$ depends on both $z=x+iy$ and its complex conjugate $z^{*}=x-iy$, it is typically not holomorphic and the usual derivative does not exist. Wirtinger calculus treats $z$ and $z^{*}$ as independent variables. For any real-valued $\mathcal{H}(z,z^{*})$, Wirtinger partial derivatives are defined as
$\partialderivative{\mathcal{H}}{z_{i}}\triangleq\frac{1}{2}\left(\partialderivative{\mathcal{H}}{x_{i}}-i\partialderivative{\mathcal{H}}{y_{i}}\right),$
$\partialderivative{\mathcal{H}}{z_{i}^{*}}\triangleq\frac{1}{2}\left(\partialderivative{\mathcal{H}}{x_{i}}+i\partialderivative{\mathcal{H}}{y_{i}}\right).$
As a consequence of the definitions above, we have
$\partialderivative{\mathcal{H}}{z_{i}}=\left(\partialderivative{\mathcal{H}}{z_{i}^{*}}\right)^{*}$
$\partialderivative{\mathcal{H}}{z_{i}^{*}}=\left(\partialderivative{\mathcal{H}}{z_{i}}\right)^{*}.$
While the formal definitions are expressed using $x$ and $y$, the Wirtinger partial derivatives behave as expected for a function defined in terms of $z$ and $z^{*}$. For example, one can easily verify that $\partialderivative{z}[zz^{*}]=z^{*}$ and $\partialderivative{z^{*}}[zz^{*}]=z$.
Additionally, a function $\mathcal{H}$ is holomorphic if and only if $\partial\mathcal{H}/\partial z_{i}^{*}=0$ for all $z_{i}$. Intuitively, Wirtinger calculus treats $\mathcal{H}$ as a function over $\mathbb{R}^{2n}$ instead of $\mathbb{C}^{n}$, using appropriately defined isomorphisms to transform between $\mathcal{H}(x,y):\mathbb{R}^{2n}\to\mathbb{R}$, $\mathcal{H}(z,z^{*}):\mathbb{C}^{2n}\to\mathbb{R}$ and $\mathcal{H}(z):\mathbb{C}^{n}\to\mathbb{R}$. From this perspective, typical objects from multivariate calculus (gradients, Hessians, $\dots$) have a natural extension to real-valued, non-holomorphic functions. Additional details can be found in the excellent introduction to the subject by Kreutz-Delgado [73]. Crucial for our purposes is the definition of the Wirtinger gradient
$\nabla_{z}\mathcal{H}\triangleq\left(\partialderivative{\mathcal{H}}{z}\right)^{\dagger},$
where $\partialderivative{\mathcal{H}}{\bm{z}}\in\mathbb{C}^{2n}$ is the vector
$\left(\partialderivative{\mathcal{H}}{z_{1}},\dots,\partialderivative{\mathcal{H}}{z_{n}},\partialderivative{\mathcal{H}}{z_{1}^{*}},\dots,\partialderivative{\mathcal{H}}{z_{n}^{*}}\right)^{\top}$
and $(\cdot)^{\dagger}$ denotes a conjugate transpose. Since Wirtinger partials have the bijective relationship $\partialderivative{\mathcal{H}}{z_{\ell}}=\left(\partialderivative{\mathcal{H}}{z_{\ell}^{*}}\right)^{*}$, we focus only on the non-conjugated components.
Example: Consider an energy function $\mathcal{H}(z,z^{*})=z\,z^{*}.$ Treating $z^{*}$ as constant when differentiating with respect to $z$ (and vice versa), we obtain
$\frac{\partial\mathcal{H}}{\partial z}=z^{*},\qquad\frac{\partial\mathcal{H}}{\partial z^{*}}=z.$
Since $\partial\mathcal{H}/\partial z^{*}=z\neq 0$, $\mathcal{H}$ is non-holomorphic.
Energy Minimization via Gradient Descent: For a real-valued Hamiltonian $\mathcal{H}(z,z^{*})$, a small update $\Delta z$ produces
$\Delta\mathcal{H}\approx\frac{\partial\mathcal{H}}{\partial z}\,\Delta z+\frac{\partial\mathcal{H}}{\partial z^{*}}\,\Delta z^{*}.$
Recall that $\partialderivative{\mathcal{H}}{z^{*}}=\left(\partialderivative{\mathcal{H}}{z}\right)^{*}$ and vice versa. Therefore, to minimize $\Delta\mathcal{H}$, it suffices to choose
$\Delta z\propto-\partialderivative{\mathcal{H}}{z^{*}}$ (35)
which produces
$\Delta\mathcal{H}\propto-2\left\lVert\partialderivative{\mathcal{H}}{z^{*}}\right\rVert^{2}.$
Choosing a “step size” of $\eta$, we therefore have a well-defined gradient descent scheme:
$z^{k+1}\leftarrow z^{k}-\eta\,\frac{\partial\mathcal{H}}{\partial z^{*}}.$
In the previous example $\partial\mathcal{H}/\partial z^{*}=z$, so iterating $z^{k+1}\leftarrow z^{k}-\eta\,z^{k}$ drives $z\to 0$, the minimizer of the energy $\mathcal{H}=|z|^{2}$.
Appendix C General PUBO Wirtinger Potentials
As defined in Sec. III.2, we can express the “complete” $k^{\textrm{th}}$ order PUBO Hamiltonian $\mathcal{H}^{C}_{k}:\mathbb{C}^{n}\to\mathbb{R}$ as
$\mathcal{H}^{C}_{k}(z)=\frac{1}{\lfloor k/2\rfloor}\sum_{p=1}^{\lfloor k/2\rfloor}\mathcal{H}^{p}_{k}(z),$
where each $p$-potential has $m_{p}=2\binom{k}{p}$ terms and each $\mathcal{H}^{p}_{k}:\mathbb{C}^{n}\to\mathbb{R}$ has the form
$\mathcal{H}^{p}_{k}(z)=\frac{1}{2^{k}}\sum_{j=1}^{m_{p}}c_{p,j}\left(y^{(1)}_{p,j}\dots y^{(p)}_{p,j}(y^{(p+1)}_{p,j})^{*}\dots(y^{(k)}_{p,j})^{*}\right$.
$+\left.(y^{(1)}_{p,j})^{*}\dots(y^{(p)}_{p,j})^{*}y^{(p+1)}_{p,j}\dots y^{(k)}_{p,j}\right)$,
where $p$ denotes the number of non-conjugated terms in the first product and each term $y^{(q)}_{p,j}$ refers to some complex variable $z_{\ell}$ appearing as the $q^{\text{th}}$ term in the $j^{\textrm{th}}$ product.